find the length of the arc on a circle of radius r=5 yard intercepted by a central angle 0=70degree..................the zero at the end has a line through it, and the 70 has a degree circle but i dont know how to make those with my keyboard..
Found 2 solutions by kos, Edwin McCravy:
Answer by kos(2)

= arc length = theta/360deg * 2pie r
= theta/360deg * 2 * 22/7 * 5
= 70deg/360 * 2 * 22/7 * 5
= 6 approximately

Answer by Edwin McCravy(20086)

find the length of the arc on a circle of radius r=5 yard intercepted by a
central angle 0=70degree..................the zero at the end has a line
through it, and the 70 has a degree circle but i dont know how to make those
with my keyboard..

The "zero with a line through it" is the Greek letter "phi" ø, which is
pronounced "fee" in mathematics (and by Greeks), but college fraternity
guys are dumb and pronounce it "figh".  :-) 

To make the "degree circle" ° on the keyboard:

1. Make sure NumLock on the far right side of your keyboard is ON.
   If it isn't ON, press the NumLock key and a small light will
   turn on to indicate that NumLock is ON.

2. Hold down the Alt key with your left forefinger.

3. While holding down the Alt key with your left forefinger, type 248
   with your right hand on the number keys at the far right side of 
   your keyboard (NOT on the number keys above the letter keys.)

4. Release the Alt key and the ° will "magically" appear on the screen.

To make the ø, follow the same 4 steps but type 0248 instead of just 248.

The length of arc is given by the formula

s = rq where r is the radius and q is the angle IN RADIANS, not degrees.

So first, change 70° to radians by multiplying 70° by p/180°

70° × p/180° = 7p/18

Then substitute r = 5 and q = 7p/18

s = (5)(7p/18) = 6.108652382 yd. approximately.

Edwin